OpenStudy (nancyg94):

Please show steps to simplify this: 8(1-(1/2)^9)/(1-(1/2)

MARC (marc_d):

\[=\frac{ 8(1-(\frac{ 1 }{ 2 })^9) }{ 1-\frac{ 1 }{ 2 } }\]like this?

MARC (marc_d):

@Nancyg94

OpenStudy (nancyg94):

Yes!

MARC (marc_d):

okay..

MARC (marc_d):

\[=\frac{ 8(1-\frac{ 1 }{ 512 }) }{ -\frac{ 1 }{ 2 } }\]

MARC (marc_d):

\[\color{red}{(\frac{ 1 }{ 2 })^9=\frac{ 1 }{ 512 }}\]

MARC (marc_d):

\[\color{blue}{1-\frac{ 1 }{ 2 }=-\frac{ 1 }{ 2 }}\]

MARC (marc_d):

\[=\frac{ 8(1-\color{red}{\frac{ 1 }{ 512 }}) }{ \color{blue}{-\frac{ 1 }{ 2 }} }\]

OpenStudy (nancyg94):

I thought 1-(1/2)= (1/2)

MARC (marc_d):

1-(1/2) is the same as 1-1/2

MARC (marc_d):

\[1\color{red}{-1}(\frac{ 1 }{ 2 })=1\color{red}{-1}\times \frac{ 1 }{ 2 }=1-\frac{ 1 }{ 2 }\]

MARC (marc_d):

\[\color{green}{1-\frac{ 1 }{ 512 }=\frac{ 511 }{ 512 }}\]\[=\frac{ 8(\color{green}{\frac{ 511 }{ 512 }}) }{ -\frac{ 1 }{ 2 } }\]

MARC (marc_d):

\[\color{aqua}{8(\frac{ 511 }{ 512 })=1\cancel{8}\times \frac{ 511 }{ 64\cancel{512} }}\]\[=\frac{ \color{aqua}{\frac{ 511 }{ 64 }} }{ -\frac{ 1 }{ 2 } }\]

MARC (marc_d):

\[=\frac{ 511 }{ 64 }\div -\frac{ 1 }{ 2 }\]\[=\frac{ 511 }{ 64 }\times-\frac{ 2 }{ 1 }\]

MARC (marc_d):

\[=\frac{ 511 }{ 32\cancel{64} }\times-\frac{ 1\cancel{2} }{ 1 }\]\[=\frac{ 511 }{ 32 }\times-1\]\[=?\]solve it! @Nancyg94

OpenStudy (nancyg94):

-511/32

OpenStudy (nancyg94):

Thanks

MARC (marc_d):

yep,that's right!

MARC (marc_d):

yw!